OpenStudy (anonymous):

Solve 4x + 3y = 13 x + 5y = 33 A. (2, 7) B. (–2, 7) C. (2, –7) D. (–2, –7)

OpenStudy (anonymous):

So now we see how this equation has to be for it to be correct. a. 8 + 21 = 13 --- NO b. -8 + 21 = 13 --- YES c. 8 - 21 = 13 --- NO d. -8 - 21 = 13 --- NO So your answer is B, just by using the first equation. If you are unsure, just use the second equation and use substitution as well on that one. You have x + 5y = 33 So it'll be... a. 2 + 35 = 33 --- NO b. -2 + 35 = 33 --- YES c. 2 - 35 = 33 --- NO d. -2 - 35 = 33 --- NO Both equations have the answer B, so for SURE YOUR ANSWER IS GOING TO BE B!

OpenStudy (anonymous):

Youre the bomb!!! :) Thankyou

OpenStudy (anonymous):

Haha You're welcome :D